# suppSupplementary Material References

### v1

@@ B.2 Sparse Matrix Multiplication
$\begin{array}[]{llllllllll}\underset{(K\times K)}{\bm{P}_{\bm{+}}^{(t)}}:=\left[\begin{array}[]{cccc}\log\dfrac{p_{11}^{(t)}(``+")}{p_{11}^{(t)}(``0")}&\log\dfrac{p_{12}^{(t)}(``+")}{p_{12}^{(t)}(``0")}&\ldots&\log\dfrac{p_{1K}^{(t)}(``+")}{p_{1K}^{(t)}(``0")}\\ \log\dfrac{p_{21}^{(t)}(``+")}{p_{21}^{(t)}(``0")}&\log\dfrac{p_{22}^{(t)}(``+")}{p_{22}^{(t)}(``0")}&\ldots&\log\dfrac{p_{2K}^{(t)}(``+")}{p_{2K}^{(t)}(``0")}\\ \vdots&\vdots&\ddots&\vdots\\ \log\dfrac{p_{K1}^{(t)}(``+")}{p_{K1}^{(t)}(``0")}&\log\dfrac{p_{K2}^{(t)}(``+")}{p_{K2}^{(t)}(``0")}&\ldots&\log\dfrac{p_{KK}^{(t)}(``+")}{p_{KK}^{(t)}(``0")}\end{array}\right]\end{array}$
@@
$\begin{array}[]{llllllllll}\underset{(K\times K)}{\bm{P}_{\bm{-}}^{(t)}}:=\left[\begin{array}[]{cccc}\log\dfrac{p_{11}^{(t)}(``-")}{p_{11}^{(t)}(``0")}&\log\dfrac{p_{12}^{(t)}(``-")}{p_{12}^{(t)}(``0")}&\ldots&\log\dfrac{p_{1K}^{(t)}(``-")}{p_{1K}^{(t)}(``0")}\\ \log\dfrac{p_{21}^{(t)}(``-")}{p_{21}^{(t)}(``0")}&\log\dfrac{p_{22}^{(t)}(``-")}{p_{22}^{(t)}(``0")}&\ldots&\log\dfrac{p_{2K}^{(t)}(``-")}{p_{2K}^{(t)}(``0")}\\ \vdots&\vdots&\ddots&\vdots\\ \log\dfrac{p_{K1}^{(t)}(``-")}{p_{K1}^{(t)}(``0")}&\log\dfrac{p_{K2}^{(t)}(``-")}{p_{K2}^{(t)}(``0")}&\ldots&\log\dfrac{p_{KK}^{(t)}(``-")}{p_{KK}^{(t)}(``0")}\end{array}\right].\end{array}$
@@ B.3 Update Rules
$\begin{array}[]{llllllllll}\bm{\alpha}^{(t+1)}&\coloneqq&\underset{\bm{\alpha}}{\operatorname*{arg\,max}}\,Q(\mbox{\boldmath$\gamma$}^{(t)},\bm{\theta}^{(t)},\bm{\alpha}^{(t)};\bm{\alpha})\vspace{0.25cm}\\ \gamma_{k}^{(t+1)}&\coloneqq&\dfrac{1}{N}\displaystyle\sum\limits_{i=1}^{N}\alpha_{i,k}^{(t+1)},\,\text{ for }k=1,\dots,K\vspace{0.25cm}\\ \pi_{k,l}^{(t+1)}(y)&\coloneqq&\dfrac{\displaystyle\sum\limits_{i<j}\alpha_{i,k}^{(t+1)}\alpha_{j,l}^{(t+1)}\mathbb{I}(y_{i,j}=y)}{\displaystyle\sum\limits_{i<j}\alpha_{i,k}^{(t+1)}\alpha_{j,l}^{(t+1)}}\end{array}$
@@ C.2 Goodness-of-Fit
The results show that while the sparse, low-degree nature of the Wikipedia network is captured by the simulations, the statistics distribution of the observed network is not replicated with high accuracy. Overall, the in-sample goodness-of-fit analysis aligns with the out-of-sample analysis, indicating that the more complex models, Partial Triad and Full Triad, best describe the data.

### v2

@@ B.2 Sparse Matrix Multiplication
$\begin{array}[]{llllllllll}\underset{(K\times K)}{\bm{P}_{\bm{+}}^{(t)}}:=\left[\begin{array}[]{cccc}\log\dfrac{p_{1,1}^{(t)}(``+")}{p_{1,1}^{(t)}(``0")}&\log\dfrac{p_{1,2}^{(t)}(``+")}{p_{1,2}^{(t)}(``0")}&\ldots&\log\dfrac{p_{1,K}^{(t)}(``+")}{p_{1,K}^{(t)}(``0")}\\ \log\dfrac{p_{2,1}^{(t)}(``+")}{p_{2,1}^{(t)}(``0")}&\log\dfrac{p_{2,2}^{(t)}(``+")}{p_{2,2}^{(t)}(``0")}&\ldots&\log\dfrac{p_{2,K}^{(t)}(``+")}{p_{2,K}^{(t)}(``0")}\\ \vdots&\vdots&\ddots&\vdots\\ \log\dfrac{p_{K,1}^{(t)}(``+")}{p_{K,1}^{(t)}(``0")}&\log\dfrac{p_{K,2}^{(t)}(``+")}{p_{K,2}^{(t)}(``0")}&\ldots&\log\dfrac{p_{K,K}^{(t)}(``+")}{p_{K,K}^{(t)}(``0")}\end{array}\right]\end{array}$
@@
$\begin{array}[]{llllllllll}\underset{(K\times K)}{\bm{P}_{\bm{-}}^{(t)}}:=\left[\begin{array}[]{cccc}\log\dfrac{p_{1,1}^{(t)}(``-")}{p_{1,1}^{(t)}(``0")}&\log\dfrac{p_{1,2}^{(t)}(``-")}{p_{1,2}^{(t)}(``0")}&\ldots&\log\dfrac{p_{1,K}^{(t)}(``-")}{p_{1,K}^{(t)}(``0")}\\ \log\dfrac{p_{2,1}^{(t)}(``-")}{p_{2,1}^{(t)}(``0")}&\log\dfrac{p_{2,2}^{(t)}(``-")}{p_{2,2}^{(t)}(``0")}&\ldots&\log\dfrac{p_{2,K}^{(t)}(``-")}{p_{2,K}^{(t)}(``0")}\\ \vdots&\vdots&\ddots&\vdots\\ \log\dfrac{p_{K,1}^{(t)}(``-")}{p_{K,1}^{(t)}(``0")}&\log\dfrac{p_{K,2}^{(t)}(``-")}{p_{K,2}^{(t)}(``0")}&\ldots&\log\dfrac{p_{K,K}^{(t)}(``-")}{p_{K,K}^{(t)}(``0")}\end{array}\right].\end{array}$
@@ B.3 Update Rules
$\begin{array}[]{llllllllll}\bm{\alpha}^{(t+1)}&\coloneqq&\underset{\bm{\alpha}}{\operatorname*{arg\,max}}\,Q(\mbox{\boldmath$\gamma$}^{(t)},\bm{\theta}^{(t)},\bm{\alpha}^{(t)};\bm{\alpha})\vspace{0.25cm}\\ \gamma_{k}^{(t+1)}&\coloneqq&\dfrac{1}{N}\displaystyle\sum\limits_{i=1}^{N}\alpha_{i,k}^{(t+1)},\,\text{ for }k=1,\dots,K\vspace{0.25cm}\\ p_{k,l}^{(t+1)}(y)&\coloneqq&\dfrac{\displaystyle\sum\limits_{i<j}\alpha_{i,k}^{(t+1)}\alpha_{j,l}^{(t+1)}\mathbb{I}(y_{i,j}=y)}{\displaystyle\sum\limits_{i<j}\alpha_{i,k}^{(t+1)}\alpha_{j,l}^{(t+1)}}\end{array}$
@@ C.2 Goodness-of-Fit
The results show that while the sparse, low-degree nature of the Wikipedia network is captured by the simulations, the statistics distribution of the observed network is not replicated with high accuracy. Overall, the in-sample goodness-of-fit analysis aligns with the out-of-sample analysis, indicating that the more complex models, “Partial Triad” and “Full Triad”, best describe the data.